HUBBLE SPACE TELESCOPE DAILY REPORT #2246 PERIOD COVERED: 0000Z (UTC) 11/06/98 - 0000Z (UTC) 11/09/98 Daily Status Report as of 313/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Eight Sets of WF/PC-2 7713 (WF/P-2 Cycle 7 Supplemental Darks pt3/3) The WF/PC-2 was used to make supplemental dark calibration images. This dark calibration program obtains three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were completed as planned, and no anomalies were noted. 1.2 Completed Six Sets of STIS/CCD 7948 (CCD Dark and Bias Monitor -- Cont. to Dec. 98) The Space Telescope Imaging Spectrograph (CCD) was used to measure CCD dark current and bias rates for the two supported amplifier settings {GAIN=1 and GAIN=4} in order to gather data for making superdarks and superbiases. The data will be used to track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. The observations were executed as scheduled, and no anomalies were noted. 1.3 Completed Eight Sets of WF/PC-2 7909 (POMS Test Proposal: WF II Parallel Archive Proposal) The WF/PC-2 was used to make parallel observations as part of the POMS test proposal. This test proposal is designed to simulate scientific plans. The observations were completed as planned, and no problems were reported. 1.4 Completed NIC/1/2/3 7227 (A Search for Low Mass/Sub-Luminous Companions to M-Stars) The Near Infrared Camera (Cameras One, Two and Three) was used to make a search for low mass/sub-luminous companions to low mass M class stars. Knowledge of stellar and sub-stellar masses and luminosities at and below the ~0.08 M{sun} H-burning limit is of fundamental importance. A dark calibration image was also taken. The observations were completed as planned, and no problems were reported. 1.5 Completed STIS/CCD/MA1/MA2 7400 (Evidence from Boron for Massive Star Evolution) The Space Telescope Imaging Telescope (CCD/MA1/MA2) was used to observe the B III and B II resonance lines at 2065.8 and 1362.5 Angstroms, respectively, in a sample of main-sequence and slightly evolved blue stars using the HST STIS MAMA detectors and echelle gratings. The proposal completed with no reported problems. 1.6 Completed Fourteen Sets of NIC/2/3 8082 (NICMOS Pure Parallel Fall 98) 30] The Near Infrared Camera (Cameras Two and Three) was used to make pure parallel observations while another instrument is prime. The observations were executed as scheduled, and no anomalies were noted. 1.7 Completed STIS/CCD/MA1 7299 (Interfaces Between Hot and Cold Gas: The Superbubble N51D in the LMC) The Space Telescope Imaging Telescope (CCD/MA1) was used to take E140M echelle spectra for four probes of the X -ray bright superbubble N 51 D in the LMC, and then to compare those spectra with WFPC2 images of the H Alpha and S II emission in this region to be taken in parallel. The observation completed without incident. 1.8 Completed WF/PC-2 8059 (POMS Test Proposal: WF-2 Targeted Parallel Archive Proposal) The WF/PC-2 was used to make observations in parallel with the prime science observations. The parallel opportunities available with WF/PC-2 in the neighborhood of bright galaxies are treated in a slightly different way from the normal pure parallels. Local group galaxies offer the opportunity for a closer look at young stellar populations. The observations were completed as planned, and no anomalies were reported. 1.9 Completed STIS/CCD/MA1/MA2 7317 (Spatial Distribution of Ozone and Sulphur Dioxide on Ganymede and Europa) The Space Telescope Imaging Telescope (CCD/MA1/MA2) was used to investigate the O_2 and O_3 in spectra of Ganymede's trailing hemisphere. Ganymede's icy surface is an unexpectedly complex and dynamic environment. Europa, exposed to a slightly different radiation environment, has SO_2 on its trailing hemisphere with no trace of either O_3 or O_2. The new capability to measurespectra at 0.1 arcsec spatial scales offered by STIS gives us the opportunity to address such questions by mapping the distribution of O_3 and OH on Ganymede and SO_2 on Europa. The proposal completed as planned, with no reported problems. 1.10 Completed WF/PC-2 6776 (A Comparative HST Imaging Study of Radio Galaxies and Hosts) A clear understanding of the nature of the host galaxies of radio-quiet quasars, radio-loud quasars and radio galaxies is of central importance in any attempt to unify and relate the three main classes of powerful active galaxies. Existing HST studies of quasar hosts, whilst producing many interesting new results, have yielded surprisingly little hard information on the properties of the hosts themselves. This imaging study using the WF/PC-2 will remedy this situation by undertaking redder, line-free observations of our three matched AGN samples. The proposal completed with no reported problems. 1.11 Completed NIC/1 7867 (NICMOS Imaging of the Dusty CfA Seyfert Nuclei) The Near Infrared Camera (Camera One) was used to make observations of the Seyfert Galaxy NGC-5347. These data will allow us to determine if the Seyfert Type 1/2 dichotomy is best explained by the unified model, the evolutionary picture, or a combination of both of these scenarios. The observations were executed as scheduled, and no anomalies were reported. 1.12 Completed NIC/1/2/3 7225 (Proper Motion of Stars Near SGR-A*) The Near Infrared Camera (Cameras One, Two and Three) was used to make astrometric proper motion observations of stars in the vicinity of Sagittarius-A (a.k.a., NGC-4147). Sgr-A may be a black hole with a mass close to a million solar masses. The observations were completed as planned, and no anomalies were reported. 1.13 Completed Two Sets of NIC/2 7226 (Search for Massive Jupiters) The Near Infrared Camera (Camera Two) was used to make a corographic search for massive planets around nearby, young main sequence stars. Because of the extreme youth of these objects, any low-mass brown dwarf and planetary companions will still be in a higher luminosity phase and thus easily detectable. A dark calibration image was also taken. The observations were executed as scheduled, and no anomalies were noted. 1.14 Completed WF/PC-2/STIS/CCD 7717 (Spectroscopy of Jupiter, Saturn, Titan and the Moon in the UV) The WF/PC-2 and the Space Telescope Imaging Spectrograph (CCD) were used to observe the Moon, Jupiter, Saturn and Titan in the UV. The goal is to determine weak absorption features, due to atmospheric gases, as well as Raman scattering features in the spectra of Jupiter, Saturn and Titan. The Moon will be used as a solar analog to reduce planetary fluxes to reflectivities. The proposal completed as scheduled, with no reported problems. 1.15 Completed Three Sets of WF/PC-2/NIC/1/2/3 7419 (The Low Mass Stellar Luminosity Function in Dwarf Spheroidal Galaxies) The WF/PC-2 and the Near Infrared Camera (Cameras One, Two and Three) was used to study of the low mass stellar luminosity function in Ursa Minor. This class of dSph satellite galaxies of the Milky Way provide the nearest, and best studied, examples of dark-matter dominated systems. They have stellar velocity dispersions well in excess of those consistent with their mass being predominantly in a stellar mass function at all like that of the solar neighborhood, or of the Galactic globular clusters. The observations were completed as planned, and no problems were reported. 1.16 Completed FOC/96 7923 (FOC F/96 Turnon) The Faint Object Camera (f/96) was used to take an internal flat and a dark to ensure that the 30-day limit for the FOC is not exceeded. The observations were executed as scheduled, and no anomalies were noted. 1.17 Completed S/C 5582 (FOC/48 Monthly Activation (Camera Section Only): Cycle 4) This proposal was executed as scheduled, and no anomalies were noted. 1.18 Completed STIS/CCD/MA1/MA2 7673 (STIS MAMA Sensitivity and Focus Monitor) The Space Telescope Imaging Spectrograph (CCD/MA1/MA2) was used to make observations of standard star AGK+81D266 in order to perform a MAMA sensitivity and focus monitoring program. This is to monitor sensitivity of each MAMA grating mode to detect any change due to contamination or other causes. The observations were executed as scheduled, and no anomalies were noted. 1.19 Completed Six Sets of STIS/CCD/MA1/MA2 7389 (Chromospheres and Transition Layers in Hyades and Pleiades F stars) The Space Telescope Imaging Spectrograph (CCD, MA1 and MA2) was used to make observations of the F class stars HD28608, HD27561, HD30738, HD28568, HD28483, and HD31845 which is located in the Hyades star cluster. This proposal seeks to obtain additional empirical constraints on the heating mechanisms for stellar chromospheres, transition layers and coronae. A wavelenth calibration obseration was also taken. The observations were completed as planned, and no problems were reported. 1.20 Completed Two Sets of NIC/1/2/3 7233 (Dust Disks Around Main Sequence Stars) The Near Infrared Camera (Camera One, Two and Three) was used to make observations of the dust disk around main sequence stars. The observations were made with the coronagraph to minimize the effects of glare from the bright central star. A dark calibration image was also taken. The observations were completed as scheduled, and no anomalies were reported. 1.21 Completed WF/PC-2 6794 (Proper Motions of Herbig-Haro Jets) The WF/PC-2 was used to make observations of the Herbig-Haro object HH111 in order to measure the proper-motions of these high Mach number flows of its two opposing jets. The tangential velocities of several hundred km/s will produce displacements between our first- and second-epoch images of 2-4 pixels after only two years. The observations were completed as planned, and no anomalies were reported. 1.22 Completed WF/PC-2 6822 (Ionized Gas With Broad Emission Lines in the Nuclei of Ellipticals) The WF/PC-2 was used to make observations of the large elliptical galaxy NGC-2325 in order to study the broad emission lines in the nucleus. Ionized gas disks in the central regions of elliptical galaxies have proven to be a powerful tool for the determination of the true spatial shape and the central mass-to-light ratio of their host galaxies. The observations were completed as planned, and no problems were reported. 1.23 Completed NIC/2 7894 (Completing A Near-Infrared Search for Very Low Mass Companions to Stars within 10 pc) The Near Infrared Camera (Camera Two) was used to make observations searching for very low mass (VLM) companions to the star G239-025. This search will be the largest, most sensitive, volume-limited search for VLM companions ever undertaken. The observations were completed as planned, and no anomalies were reported. 1.24 Completed NIC/1/2/3 7902 (Photometric Monitoring, Photometric Monitoring-II) The Near Infrared Camera (Cameras One, Two and Three) was used to make observations of absolute standard, solar analog star P330E-CAL as part of a photometric monitoring program. Observations were made at several different wavelengths spanning the NICMOS wavelength range in each camera NIC/1, NIC/2, and NIC/3 to monitor the photometric stability of the NICMOS cameras as a function of wavelength. The observations were executed as scheduled, and no anomalies were noted. 1.25 Completed WF/PC-2 6666 (The Mass Function and Binary Star Fraction of the Trapezium Cluster) The WF/PC-2 was used to take deep images of the Trapezium cluster to determine the low-mass end of the initial mass function, to search for brown dwarfs, and to explore the close binary population in the cluster. The observations completed as planned with no problems. 1.26 Completed Nine Sets of WF/PC-2 8061 (POMS Test Proposal: WF-II Backup Parallel Archive Proposal) This is a POMS test proposal designed to simulate scientific plans. The proposal was executed as planned, and no anomalies were reported. 1.27 Completed WF/PC-2 7377 (Dwarf Elliptical Galaxy Snapshot Survey II) The WF/PC-2 was used to make observations of the dwarf elliptical galaxy FCC282 as part of Snapshot Survey II. This survey will provide important information on the globular cluster systems of such galaxies and the properties of their cores and nuclei. The observations were completed as planned, and no anomalies were reported. 1.28 Completed NIC/1/2/3 6482 (Weak Lensing in the Field of Luminous Quasars) The Near Infrared Camera (Cameras One, Two, and Three) was used to observe some relatively bright QSOs which have long been suspected to be gravitationally magnified. We have found along the line of sight towards some of them a coherent shear pattern of faint distant galaxies. Moreover there appear to be overdensities of galaxies concentric with the shear patterns which are probably responsible for the weak lensing effect and the QSO magnifications. The observations completed as planned, with no reported anomalies. 1.29 Completed NIC/1/2/3 7247 (Environments of T-Tauri Stars) The Near Infrared Camera (Cameras One, Two and Three) was used to make observations of the T-Tauri class star HL-Tau. T-Tauri stars are low {0.5-2 solar} mass, pre-main-sequence stars. Many may have circumstellar disks - theoretical models involving accretion disks have successfully explained some of the basic observational properties of these young objects. The observations were executed as scheduled, and no anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 29 Successful : 29 Scheduled Reacquisitions : 14 Successful : 11 Following a successful acquisition at 313/0050Z, reacquisitions at 313/0210Z, at 313/0345Z, and at 313/0520Z failed due to a walkdown failure. NICMOS (7963, 7226, and 7227), STIS (7950), and WF/PC-2 (7620 and 7713) were affected. 2.2 FHST Updates: Scheduled : 67 Successful : 67 2.3 Operations Notes: The Battery #2 capacity check test completed nominally at 31-/2005Z. Gyro re-configuration (gyros 3, 1, 5, and 2) was completed nominally at 310/22201Z. An attitude reference update of 650 arcseconds of error, resulting from the re-configuration, was uplinked at 310/2240Z as planned. During the first T2 slew of 100 degrees, approximately .75 arcseconds of slew attitude error was determined. The first FGS acquisition following the re-configuration, at 311/0030Z, was successful. Losses of lock did occur during the planned gyro mode switch to high mode to obtain high mode bias using the observer. The science timeline was intercepted at 311/0243Z following a T2 slew and successful updates. Using SR-1, the SSR EDAC error counter was cleared eight times. At 310/234730Z, 599 correctable EDAC errors occurred during a playback. The same number of errors occurrd during the re-playback. HSTAR 6759 was written. At 311/133455Z, per HSTAR 6760, there was a NICMOS single bit error. Per HSTAR 6762, there were NICMOS single bit errors at 312/012856Z, at 312/013255Z, and at 312/015256Z. Then, as described in HSTAR 6763, another NICMOS single bit error occurred at 312/170857Z. In each case, the NICMOS internal log was dumped, and the ground system limit was modified. ESTR-3 reconditioning was performed at 312/0210Z per ROP RD-7. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH